# When Attention Collapses: How Degenerate Layers in LLMs Enable Smaller, Stronger Models

### v3

@@ Appendix C Supplementary Experiments
Figure 16: A 12-Layer GPT-2 medium variant derived using Inheritune converges faster and generalizes better than a same-sized models trained from scratch and with Logit-based distillation with teacher initialization baseline. Three 12-layer GPT-2 medium variants were trained: (1) a distilled model initialized with alternate layers from a 24-layer GPT-2 medium teacher, following the DistillBERT setup Sanh et al. (2019); (2) a model trained from scratch (vanilla training); and (3) a model trained using the $\mathsf{Inheritune}$ recipe. The model trained with $\mathsf{Inheritune}$ outperforms both the distillation-based model and the one trained from scratch, demonstrating the effectiveness of our approach.
(a) Vanilla GPT2-Medium 24 layers
(b) Our GPT2-Medium 16 layer variant trained with $\mathsf{Inheritune}$
Figure 17: Rank collapse in deeper layers and its mitigation through $\mathsf{Inheritune}$. The maximum (max) rank across all attention heads for each layer is plotted, following the methodology in Fig. 1 (a) Analysis of a 24-layer GPT2 medium model reveals rank-1 attention matrices in later layers (those beyond the halfway point), indicating rank collapse. Specifically, 3 out of the last 12 later layers exhibit rank-1 attention matrices (mean rank accross all the 100 runs). (b) Our 16-layer GPT2 medium variant, trained with $\mathsf{Inheritune}$, demonstrates improved rank across all layers, highlighting the effectiveness of our approach. Notably, none of the later layers in our 16-layer variant exhibit rank-1 attention matrices.
(a) Vanilla GPT2-xLarge 48 layers
(b) Our GPT2-xLarge 24 layer variant trained with $\mathsf{Inheritune}$
Figure 18: Rank collapse worsens for larger LLMs, $\mathsf{Inheritune}$ helps to mitigate rank collapse. The maximum (max) rank across all attention heads for each layer is plotted, following the methodology in Fig. 1 (a) Analysis of a 48-layer GPT2 xLarge model reveals rank-1 attention matrices in later layers (those beyond the halfway point), indicating rank collapse. Specifically, 22 out of the last 24 later layers exhibit rank-1 attention matrices (mean rank across all the 100 runs). (b) Our 24-layer GPT2 xLarge variant, trained with $\mathsf{Inheritune}$, demonstrates improved rank across all layers, highlighting the effectiveness of our approach. Notably, 2 out of 12 of the later layers in our 24-layer variant exhibit rank-1 attention matrices.
Appendix D Extended Discussion
D.1 Discussion about Attention Sink
The term "attention sink" Xiao et al. (2024) refers to the phenomenon where the first token in a sequence receives disproportionately high attention scores compared to other tokens in the attention maps. While there is some connection with $\mathsf{Inheritune}$, as we have also observed that many attention matrices are not only rank-1 but also single-column (with all attention scores concentrated on the first token), this connection has not been explicitly established in Xiao et al. (2024) with respect to rank-1 behavior or poor training of later layers.
In contrast, as illustrated in Figure 1, we compute the maximum rank of all attention matrices within a layer. For instance, consider a layer where only 2 out of 5 attention heads exhibit attention sink behavior. This does not make the layer lazy, as attention is computed as a concatenation of activations across all heads. A lazy layer, however, has all 5 out of 5 attention heads fully degenerated, with their attention matrices being rank-1. We have provided evidence that such lazy layers are indicative of poorly trained layers.
Layers Initialization Avg max ranks Val Loss ($\downarrow$)
4 rand n/a 3.25
4 1-4 layers from vanilla GPT2 8.40 3.22
4 5-8 layers from vanilla GPT2 9.48 3.19
4 9-12 layers (lazy layers) from GPT2 1.22 3.23
Table 4: Impact of initialization strategies on GPT2-small variants. We analyzed the rank characteristics of a vanilla GPT2-small model (125M, 12 layers) trained on OpenWebText for 100K steps. Four-layer GPT2-small variants were initialized using the first 4 layers [1–4], middle 4 layers [5–8], last 4 layers [9–12], or with random initialization, and then trained for 100K steps on OpenWebText. Models initialized with the last 4 layers performed similarly to random initialization, while those initialized with layers exhibiting higher average max ranks achieved the best validation loss, regardless of proximity to the embedding layer. The training plots and rank analysis are provided in Figure 2.
Models Recipe Layers ARCE PIQA SciQ Hellaswag Lambada Average
(acc) (acc) (acc) (acc norm) (acc)
GPT-2 Medium rand init 24 51.05 61.81 74.8 30.79 20.28 47.74
rand init 16 49.92 61.92 73.3 29.56 19.54 46.84
Ours 16 51.26 61.81 73.8 30.55 23 48.08
GPT-2 Large rand init 32 52.48 64.58 75.3 32.65 22.2 49.44
rand init 16 50.34 63.11 75 30.86 21.56 48.17
Ours 16 52.9 63.55 76.1 32.14 24.06 49.75
Table 5: Models trained with Inheritune outperforms both their larger and same-size counterparts trained from scratch on average zero-shot downstream performance. For evaluation we choose accuracy (acc) and normalized accuracy (acc norm) metrics following Open LLM leaderboard Beeching et al. (2023). All the models are trained with FineWeb_edu.
Appendix E Developing a 1.5B Small Base LM in a Low Data Regime with Inheritune
Figure 19: Performance of our 1.5B base LM derived using 1B data with $\mathsf{Inheritune}$ on an average of 9 different datasets (left) and MMLU benchmark (right) that evaluates commonsense, truthfulness, natural language inference and language understanding. We compare our model’s performance with reference model-OpenLLamA-3B (2x size), other small base LMs of size 1B-2B parameters such as MPT-1.3B, OPT-1.3B, Pythia-1.4B (pre-trained from scratch) and ShearLLaMA-1.5B (pruned and continually trained using existing large base LM).
In this section, we aim to investigate the efficacy of $\mathsf{Inheritune}$ in a data and compute-constrained setting. We train a 1.5B parameter small base LM with only 1B tokens using a 3B parameter base LM on a single GPU (A6000) for less than half a day.
We assume the existence of a pre-trained reference model $\mathcal{M}_{\text{ref}}$, comprising $k$ layers, represented by $\mathsf{W}_{\text{ref}}=\{\mathsf{W}_{0},\mathsf{W}_{1},\ldots,\mathsf{W}_{k-1}\}$ trained with $\mathcal{D}_{\text{train}}$. However, this full training data is unavailable, and we only have a random tiny subset $\hat{\mathcal{D}}_{\text{train}}\sim\mathcal{D}_{\text{train}}$. We use OpenLLaMA-3B version 1 as the reference model pre-trained with 1T tokens from the RedPajama V1 dataset, which contains data from various domains such as common crawl, C4, Wikipedia, books, arXiv papers, GitHub, and Stack Exchange. We take 1B randomly sampled tokens [^3] https://huggingface.co/datasets/togethercomputer/RedPajama-Data-1T-Sample from the RedPajama dataset.
Training recipe.
To adapt $\mathsf{Inheritune}$ for this new setting, we perform step 1 and step 2 in Algorithm 1 without growing the model (i.e., we skip step 3). We use the first $n=13$ layers from our $k=26$ layer reference model. We call our small base LM Ours-1.5B(#tokens). We train our model with data repetition for eight epochs (each epoch uses all the 1B tokens) with a batch size of 131K tokens per batch. We use 1 A6000 GPU for less than half a day of training. The choice of training epochs is based on the analysis provided later in this paper (refer to Figure 22). We use the lit-gpt framework for training all small base LMs discussed in this paper. Further discussions on the training hyper-parameters can be found in the next Section.
Baseline models and evaluation.
We choose similarly sized (1-2B parameter) small base LMs trained with the RedPajama dataset and the reference base LM as primary baselines, as the quality of the pre-training data plays a key role in model development. We also include models OPT-1.3B Zhang et al. (2022) and Pythia-1.3B Biderman et al. (2023) as these models are pre-trained with a dataset similar to the RedPajama dataset. Table 7 lists the baseline models with their pre-training data.
In this study, we use few-shot accuracy, particularly 0-shot and 5-shot accuracy, on ten different downstream tasks to measure the quality of our 1.5B base LM. This evaluation of pre-trained base LLMs has been done in several prior works. Our evaluation methodology categorizes downstream tasks across four distinct genres: commonsense reasoning, natural language understanding, factuality, and natural language inference. We perform 0-shot evaluation for PIQA Bisk et al. (2020), BOOLQ Clark et al. (2019), WINOGRANDE Sakaguchi et al. (2020), WINOGRAD Kocijan et al. (2020), LOGIQA Liu et al. (2020), TruthfulQA Lin et al. (2022), MNLI Bowman et al. (2015), QNLI Wang et al. (2018) and WNLI Wang et al. (2018) datasets. Next, we perform a 5-shot evaluation on the massive multitask language understanding benchmark (MMLU) Hendrycks et al. (2020). We use the lm eval harness framework Gao et al. (2024) for the entire evaluation.
Model Commonsense Reasoning
Name (# train tokens) Reference Winograd PIQA Boolq WinoGrande Logiqa
OpenLLaMA-3B (1T) n/a 63.46 74.97 67.18 62.27 28.4
OPT-1.3B (300B) n/a 38.46 71.82 57.83 59.51 27.04
Pythia-1.4B (300B) n/a 36.54 70.89 63.12 56.99 27.65
MPT-1.3B (200B) n/a 63.46 71.44 50.89 58.09 28.26
Sheared LLaMA-1.3B (50B) LLaMA2-7B 36.54 73.45 62.02 58.17 27.34
Ours-1.5B (1B) OpenLLaMA-3B 50.96 56.47 61.68 51.69 25.19
Model Lang. Understanding & Inference Factuality
Name ( # train tokens) Reference MMLU(5) WNLI QNLI MNLI TruthfulQA
OpenLLaMA-3B (1T) n/a 27.21 50.7 51.3 37.3 35
OPT-1.3B (300B) n/a 24.96 42.25 51.29 35.82 38.67
Pythia-1.4B (300B) n/a 25.56 53.52 49.48 32.76 38.66
MPT-1.3B (200B) n/a 25.82 40.85 50.52 35.93 38.68
Sheared LLaMA-1.3B (50B) LLaMA2-7B 25.71 49.3 50.98 37.94 37.14
Ours-1.5B (1B) OpenLLaMA-3B 25.67 43.66 49.41 34.42 48.61
Table 6: Our 1.5B model achieves performance comparable to baseline models despite being trained with fewer tokens. Comparison of our target model ($\mathcal{M}_{\text{tgt}}$) derived using $\mathsf{Inheritune}$ with the reference model ($\mathcal{M}_{\text{ref}}$) and other baseline models of similar size when pre-trained from scratch and pre-trained with inherited weights and pruning. Although trained with fewer tokens, our model achieves performance comparable to the baseline models. We have highlighted all the scores in bold where our 1.5B model achieves at least 90% of the score compared to the reference LM or outperforms at least two of the baseline similar-size LMs. All the tasks are evaluated using 0-shot except MMLU, which is 5-shot. The models against which n/a is mentioned are trained from scratch.
E.1 Main Results in Low Data Regime
Table 6 presents a detailed performance evaluation across various tasks. Our 1.5B model, developed using $\mathsf{Inheritune}$, excels in 7 out of 10 individual tasks. It achieves a score of 90% or higher compared to the reference language model, which is twice its size and trained with 1000 times more data, or it outperforms at least two other base LMs of similar size trained with 50-300 times more data. Favorable scores are highlighted in bold.
Next, we compare our small LM with the MPT-1.3B [^4] https://huggingface.co/mosaicml/mpt-1b-redpajama-200b model trained from scratch with 200B tokens of RedPajama dataset and find that we match 97% accuracy in all nine downstream tasks and the MMLU (5-shot) score. Additionally, we compare with OPT-1.3B and Pythia-1.3B models, showing that we outperform both in the MMLU (5-shot) score and perform comparably on the other nine datasets. This comparison illustrates that having a large reference base LM and a subset of its pre-training data allows the inherited target size base LM to be trained remarkably more sample-efficiently than training from scratch. Extended discussions on comparisons with the ShearedLLaMa model, generated by pruning and continual training from LLaMA2-7B, are provided in the supplementary materials.
Ablation of Inheritune Across Different Model Sizes with 1B Tokens.
In the previous section, we considered a single choice of $n=k/2$, i.e., half the layers, for the size of the smaller model. Here, we investigate $\mathsf{Inheritune}$ with different choices of $n$, but the same 1B token dataset). All models use OpenLLAMA-3B as the large pre-trained reference model, with consistent training hyperparameters, changing only the choice of $n$.
We developed eight different submodels with $n=\{4,6,8,10,13,16,18,20\}$. Figure 20 shows the MMLU (5-shot) score as a function of $n$. As expected, the trend line is positive-sloping. The submodel with 20 layers slightly decreases performance, potentially due to data overfitting as the model size increases. The training details for all these submodels are consistent with the target 1.5B small base LM and are detailed in the appendix. A more comprehensive investigation on the choice of $n$—including varying both $n$ and the number of training tokens jointly and evaluating a broader set of tasks is left for future work.
Figure 20: Inheritune scales across multiple different model sizes. Utilizing the OpenLLaMA-3B as a reference large base LM, demonstrates that multiple performant small base LMs of target size can be crafted using $\mathsf{Inheritune}$ with just 1B training tokens. The MMLU (5-shot) as a function of the number of submodels.
E.2 Additional analysis with larger reference LMs and 50B data
We further analyze $\mathsf{Inheritune}$ to see the impact of it’s performance when more tokens are available. Initially for the main results we limited ourselves to 1B (i.e. 0.1%) tokens from the 1T pre-training data, here we use a 50B subset (i.e. 5%) of the pre-train data. Moreover we also extend this study to include larger base LMs of 7B parameters as reference models, employing OpenLLaMA-7B and LLaMA2-7B as reference models. For the purpose of this study we do not repeat the tokens from our 50B subset. As shown in Figure 21, we observe that there is clear improvement in overall MMLU (5-shot) score with more data. Additionally it is interesting to see that 1.5B (or 1.6B models) developed with $\mathsf{Inheritune}$ using larger reference models show even greater improvements when fed with 50B subset of non repetitive data (i.e fresh tokens). We present a Table 9 using Figure 21 to show the best MMLU (5-shot) scores achieved using different reference LMs. For developing our small base LMs using larger reference LMs we use $n$=7 (i.e. 7 layers). The training details are discussed in the following section.
Model Training Data (# tokens)
OpenLLaMA-3B v1(ref) RedPajama(1T)
Ours-1.5B* RedPajama (1B)
Shear-LLaMA-1.3B* RedPajama(50B)
MPT-1.3B RedPajama(200B)
Pythia-1.4B The Pile(300B)
OPT-1.3B Custom data(300B)
Table 7: Comparison of training data across baseline models. Overview of reference and baseline models, including their pre-training datasets and the number of tokens used during training. Note the significant variation in training data size, ranging from 1B to 1T tokens.
Models (# train tokens) GPU Count GPU Type Time (# days)
MPT-1.3B (200B) 440 A100 half
Pythia-1.4B (300B) 64 A100 4.6
TinyLLaMA-1.1B (3T) 16 A100 90
OPT-1.3B (300B) 992 A100
Sheared LLaMA-1.3B (50B) 16 A100
OpenLLaMA-3B (1T) 256 TPU v4 10
Our-1.5B (1B) 1 A6000 ~half
Table 8: Computational efficiency of Inheritune versus baseline models. Comparison of pre-training compute requirements for publicly available small base LMs and our $\mathsf{Inheritune}$-derived model. Metrics include GPU count, GPU type, and training duration, highlighting Inheritune’s significant reduction in computational resources.
Figure 21: Impact of reference model choice on Inheritune performance. MMLU (5-shot) scores for 1.5B base LMs derived using $\mathsf{Inheritune}$, trained on 50B unique tokens. Comparison across three reference models: OpenLLaMA-7B, LLaMA2-7B, and OpenLLaMA-3B. Results demonstrate Inheritune’s effectiveness with various large language models as references.
Figure 22: Performance of our 1.5B base LM derived using $\mathsf{Inheritune}$ based on existing OpenLLaMA-3B base model. Here we use 1B tokens and perform data repetition (epochs) during training. We further evaluate our model on an average of 9 different datasets (left) and MMLU benchmark (right).
Model (# tokens), ref MMLU(5)
Ours-1.6B (1B), LLaMA2-7B 24.27
Ours-1.5B (1B), OpenLLaMA-3B 25.67
Ours-1.5B (50B), OpenLLaMA-3B 25.71
Ours-1.6B (50B), LLaMA2-7B 26.07
Ours-1.6B (50B), OpenLLaMA-7B 26.72
Table 9: Performance comparison of models on the MMLU (5-shot) task. Our models, even when trained with fewer tokens, show competitive performance compared to benchmarks. We have highlighted the best MMLU 5-shot score in bold.
Model (# tokens) Data type MMLU (5-shot)
Ours-1.5B (1B) 10 epochs 24.95
Ours-1.5B (50B) 10B fresh 23.62
Ours-1.5B (1B) 20 epochs 25.46
Ours-1.5B (50B) 20B fresh 24.96
Table 10: MMLU (5-shot) scores of Our-1.5B small base LM derived using 1B data for multiple data repetition–10 epochs and 20 epochs compared to the same model trained without data repetition for 10B and 20B fresh tokens. We derive all the variants of Our-1.5B small base using $\mathsf{Inheritune}$ with OpenLLaMA-3B as reference model. The models featured in this table correspond to those discussed in Figures 21 and 22.
Ablations with number of epochs.
We ran ablations (refer Figure 22) to choose the total number of epochs (multiple passes over the data) and observe that repetition when training our 1.5B (or 1.6B) LM is helpful particularly for MMLU. We also observe that the for an average of all the 9 other datasets (i.e. except MMLU) peaks it’s performance at 5 epochs and then deteriorates. Some prior works have studied this phenomenon that the scaling of downstream tasks with data is not always linear Biderman et al. (2023).
To repeat or not to repeat the tokens.
Next we tackle the question – whether one should re-use 1B tokens for multiple epochs or use the same number of fresh tokens? Some prior works have recommended that if you have a reasonably large size dataset one can repeat it upto 4 epochs Muennighoff et al. (2023). In our study we observe that one can safely re-use 1B tokens upto 10-20 epochs as shown in Table 10. We emphasis that this phenomenon needs a through investigation in itself and we defer this to future work. The models discussed in Table are saved checkpoints during a single training run and not the final model unless otherwise specified.
E.3 Implications of Low Data Regime
In this section, we discuss some of the key implications of our work in low data regime.
Cheap and easy development of small base LMs.
Pre-training a small base LM of 1-2B parameters from scratch is extremely expensive. For instance mpt-1.3B base LM is pre-trained with 440 A100 GPUs for half a day, while the Pythia-1.4B base LM Biderman et al. (2023) utilized 64 A100-40GB GPUs for 4.6 days. Similarly, TinyLLaMA-1.1B model Peiyuan Zhang & Lu (2023) was pre-trained using 16 A100 GPUs for 3 months. Our 1.5B (1B data variant) LM shows competitive performance despite being trained with 1 A6000 GPU for less than 12 hours. The computational details are provided in Table 8, comparing the training resources of the baseline models listed in this paper. Typically small base LMs are finetuned for a specific task before deployment and are not used in it’s base form. With $\mathsf{Inheritune}$ we present a really easy and cheap way for developing a small base LM to be later finetuned before deployment.
Naive baseline for pre-training a scaled down variant of large base LMs.
Typically small variants of large base LMs are pre-trained using the same pre-training data Peiyuan Zhang & Lu (2023); Groeneveld et al. (2024). Our recipe introduces a new perspective of identifying sufficient depth without losing any generalization on the held out validation set. Next, we also show that even with a small fraction of pre-train data (randomly sampled) and few initial layers of the large base LM one can develop a small base LM. Therefore our $\mathsf{Inheritune}$ recipe has the potential to become the naive baseline for any pre-training pipeline aiming to develop a smaller variant of a large base LM.
Appendix F Implementation Details
F.1 Training details of GPT-2 models
For our main experiments, we focused on three sizes of GPT-2 models Radford et al. (2019): the vanilla GPT-2 xlarge with 1.5B parameters, GPT-2 large with 770M parameters and the vanilla GPT-2 medium with 355M parameters. We developed several variants of these models by adjusting the number of layers and hidden size. We trained all GPT-2 models with data repetition while using OpenWebText dataset, the trainset has 9B tokens and the validation set has 4.4M tokens. The key architectural configurations for the reference models, our models, and baseline models discussed in this paper are summarized in Table 11.
For all training runs, we used GELU activations, disabled bias terms, and removed dropout, following the nanoGPT codebase and Liu et al. (2023). We employed the AdamW optimizer with $\beta_{1}=0.90$ and $\beta_{2}=0.95$. The GPT-2 models were trained on a single node with 3 A100 GPUs (each with 40 GB of memory) using distributed data parallelism and gradient accumulation. In line with Liu et al. (2023), we scaled the attention logits inversely to the layer index across all GPT-2 models. Most hyperparameters were adapted from Liu et al. (2023), with key details provided below.
@@ F.1 Training details of GPT-2 models
  • Batch size: 50K tokens
  • Learning rate: $3\times 10^{-4}$,
  • Warmup steps: 2K,
  • Scheduler type: cosine decay to $\frac{1}{10}$ of max learning rate,
  • Weight decay: 0.1,
  • Gradient clipping value: 1,
@@
  • Batch size: 16K tokens
  • Learning rate: $2\times 10^{-4}$,
  • Warmup steps: 2K,
  • Scheduler type: cosine decayed to $1\times 10^{-5}$,
  • Weight decay: 0.1,
  • Gradient clipping value: 1,
@@
  • Batch size: 16K tokens
  • Learning rate: $1.5\times 10^{-4}$,
  • Warmup steps: 2K,
  • Scheduler type: cosine decayed to $1\times 10^{-5}$,
  • Weight decay: 0.1,
  • Gradient clipping value: 1,

### v4

@@ Appendix E Supplementary Experiments
Figure 20: Training curves for the 24-layer GPT-2 Medium model (full model) and three rounds of training following Inheritune recipe (to grow the model). We present the training trajectories for all GPT-2 Medium variants trained using the $\mathsf{Inheritune}$ recipe. The final model obtained after the third round ($R=3$) with $L=16$ layers matches the final validation loss of the full model. All models are trained for 100K steps on the OpenWebText dataset.
Appendix F Architectural and Training Details
F.1 GPT-2 Model configurations
For our main experiments, we focus on three sizes of GPT-2 models Radford et al. (2019): GPT-2 XLarge with 1.5B parameters, GPT-2 Large with 770M parameters, and GPT-2 Medium with 355M parameters. We developed several variants of these models by adjusting the number of layers, i.e., reducing the depth for vanilla models for to be trained with $\mathsf{Inheritune}$ and baseline mathods. In one baseline namely, the half-width variant we modified both the hidden size (and consequently, the number of attention heads) in addition to the depth, as shown in Figure 2. The key architectural configurations of the reference, proposed, and baseline models discussed in this paper are summarized in Table 6.
Model Family Type Layers Hidden Size Heads Notes
GPT-2 XLarge (1.5B) Reference 48 1600 25 Original architecture
Variants 24 1600 25 Reduced-depth
GPT-2 Large (770M) Reference 36 1280 20 Original architecture
Variants 18 1280 20 Reduced depth
GPT-2 Large (668M) Reference 32 1280 20 Custom architecture
Variants 16 1280 20 Reduced depth
GPT-2 Medium (355M) Reference 24 1024 16 Original architecture
Variants 16 1024 16 Reduced depth
GPT-2 Small (125M) Reference 12 768 12 Original architecture
Variants 4 768 12 Reduced depth
Table 6: Overview of all GPT-2 models used in this study and their architectural configurations. GPT-2 models are Pre-LN based architectures. The model configurations employed for the stacking and hybrid stacking baselines are identical to those of our variants. For the half-width baseline, we used GPT-2 variants with half the hidden size and number of attention heads.
F.2 Training details of GPT-2 models
All GPT-2 models used in this study (unless otherwise stated) were pre-trained on the OpenWebText dataset, which contains approximately 10B tokens. We employed a dataloader that samples tokens with replacement, meaning that the tokens used for training are not necessarily unique, following the approach of Liu et al. (2023). For evaluating the pre-trained models, we used the validation split of the same dataset, which contains 4.4M tokens. The sole exception to this setup is the GPT-2 models trained on the FineWeb edu with 10B tokens (Figure 7), where we used unique tokens for training by employing a dataloader where we sample without replacement.
We employed the AdamW optimizer with $\beta_{1}=0.90$ and $\beta_{2}=0.95$. All GPT-2 models were trained on a single NVIDIA A100 GPU (40 GB memory) with gradient accumulation. For the GPT-2 XLarge and its variants, we utilized an NVIDIA H100 GPU. Most hyperparameters were adapted from Liu et al. (2023), with key details discussed in this section.
@@ F.2 Training details of GPT-2 models
  • Batch size: 394K tokens
  • Learning rate: $3\times 10^{-4}$
  • Warmup steps: 2K
  • Scheduler type: cosine decayed to $1\times 10^{-5}$
  • Weight decay: 0.1
  • Gradient clipping value: 1
@@
  • Batch size: 128K tokens
  • Learning rate: $2\times 10^{-4}$
  • Warmup steps: 2K
  • Scheduler type: cosine decayed to $1\times 10^{-5}$
  • Weight decay: 0.1
  • Gradient clipping value: 1
@@
  • Batch size: 128K tokens
  • Learning rate: $1.5\times 10^{-4}$
  • Warmup steps: 2K
  • Scheduler type: cosine decayed to $1\times 10^{-5}$
  • Weight decay: 0.1
  • Gradient clipping value: 1
@@
$\text{Total\_loss}=\alpha\cdot\text{student\_loss}+(1-\alpha)\cdot\text{distillation\_loss}$
  • Model: 16-layer GPT-2 Medium variants
  • $\alpha$: 0.6
  • Batch size: 394K tokens
  • Learning rate: $3\times 10^{-4}$
  • Warmup steps: 2K
  • Scheduler type: cosine decay to $\frac{1}{10}$ of max learning rate
  • Weight decay: 0.1
  • Gradient clipping value: 1
  • Total training steps: 50K
Appendix G Additional Experiments and Discussions
G.1 Additional Downstream Evaluation
We extend our evaluation to models trained on OpenWebText, as described in Section 4 (see Table 7). For this analysis, we focus on the largest model considered in this work, GPT-2 XLarge, along with its variants. Downstream evaluation is performed on all datasets listed in Section 4, with the addition of Winogrande (Sakaguchi et al., 2020), BoolQ (Clark et al., 2019), and Wikitext (Merity et al., 2016). Note that the LAMBADA dataset is evaluated under two settings: missing-word prediction (accuracy) and language modeling (perplexity), and is therefore reported twice. All evaluations are implemented using the widely adopted lm-eval-harness.
Task Full model Ours
Accuracy-based tasks ( $\uparrow$ )
ARC-E 50.38 $\pm$ 1.03 51.22 $\pm$ 1.03
PIQA 66.70 $\pm$ 1.10 66.87 $\pm$ 1.10
SciQ 77.00 $\pm$ 1.33 79.20 $\pm$ 1.28
HellaSwag 33.65 $\pm$ 0.47 34.20 $\pm$ 0.47
LAMBADA 39.90 $\pm$ 0.68 43.30 $\pm$ 0.69
WinoGrande 51.93 $\pm$ 1.40 53.28 $\pm$ 1.40
BoolQ 57.86 $\pm$ 0.86 60.40 $\pm$ 0.86
Average 53.92 55.50
Perplexity-based tasks ( $\downarrow$ )
Wikitext 25.46 25.52
LAMBADA 20.24 16.51
Average 22.85 21.01
Table 7: Downstream evaluation of GPT-2 XLarge (1.5B) trained from scratch vs. a 24-layer model trained with Inheritune (Ours). We evaluate both models on 7 accuracy-based tasks (higher is better) and 2 perplexity-based tasks (lower is better). All models are trained on OpenWebText. Despite using half the depth, the $\mathsf{Inheritune}$ model performs favorably compared to the full-sized counterpart. Best average scores are highlighted in bold.
G.2 Additional Evidence of Attention Collapse in Modern LLMs
Following the discussion in Section B.1, we extend our analysis to several additional open-weight base LLMs (see Figures 21 and 22), including Falcon-7B [^3] https://huggingface.co/tiiuae/falcon-7b, OLMo-1B [^4] https://huggingface.co/amd/AMD-OLMo, Cerebras-GPT-2.7B [^5] https://huggingface.co/cerebras/Cerebras-GPT-2.7B, and LLaMA-3-3B [^6] https://huggingface.co/meta-llama/Llama-3.2-3B.
Many of these models incorporate modern architectural components, including Grouped Query Attention (GQA), RoPE positional embeddings, and RMS normalization applied both before and after the attention module and trained with billions-trillions of tokens. We visualized the attention heads of each model using heatmaps of head index versus layer index. Across all models, we observe a predominance of near rank-1 attention heads, indicating widespread attention collapse.
Figure 21: Rank analysis of the Falcon-7B model reveals widespread attention collapse. We analyzed Falcon-7B (version 1), which comprises 31 layers with 71 attention heads per layer, using the rank metric defined in Section 2. The results are visualized as a heatmap with layer index on one axis and head index on the other. Potent (non-collapsed) heads are shown in varying shades of green, with higher intensity indicating higher rank, while rank-collapsed (near rank-1) heads are highlighted in red. Overall, approximately 75% of attention heads exhibit rank collapse, indicating substantial degeneracy across the model.
(a) OLMo 1B
(b) Cerebras-GPT 2.7B
(c) LLaMA-3 3B
Figure 22: Rank analysis of large open-weight LLMs reveals widespread attention collapse. We analyze three base open-weight language models using the rank metric defined in Section 2. (a) OLMo (1B) exhibits rank collapse in approximately 58% of attention heads; (b) Cerebras-GPT (2.7B) shows a similar level of collapse at roughly 58%; and (c) LLaMA-3 (3B) exhibits rank collapse in about 50% of attention heads. Each heatmap plots attention head index versus layer index. Potent (non-collapsed) heads are shown in varying shades of green, with higher intensity indicating higher rank, while rank-collapsed (near rank-1) heads are highlighted in red.
G.3 Additional Results on the Functional Ineffectiveness of Lazy Layers
Following the discussion in Section 2.1 on the poor transferability of lazy layers, we conduct an additional experiment. We first train a 24-layer GPT-2 Medium model (reference) on OpenWebText for 100K steps. We then train a second model of identical size using the same data and training configuration. In this model, the deeper layers (layers 16–24) are initialized with the corresponding weights from the reference model, while the remaining layers are randomly initialized. As shown in Figure 23, reusing lazy layers consistently degrades performance relative to training all layers from scratch. This result complements our findings in Section 2.1 and provides further evidence that lazy layers do not encode transferable or reusable representations suitable for initialization. We used a batch size of 50K tokens with 2 H100 GPUs; all other training details remain unchanged (see Section F.2)
G.4 Additional Training Results in the Extended Training Regime
In this section we present a supplementary result where we have extended the training steps from 100K (as discussed in Section 4) to 200K steps to gauge the potential $\mathsf{Inheritune}$ holds for longer training runs. As shown in Figure 24 a GPT-2 medium (24 layer full model) is compared against a 18 layer GPT-2 medium variant trained following $\mathsf{Inheritune}$ recipe. We have inherited $l=18$ layers based on the previously known best configuration for GPT-2 medium variant where a 16 layer model matches the performance its full sized counterpart. We observe that the $\mathsf{Inheritune}$ model with 18 layers achieves a lower final validation loss (2.86) than the 24-layer baseline (2.87), demonstrating that $\mathsf{Inheritune}$ continues to provide benefits even with extended training. We used a batch size of 50K tokens with 2 H100 GPUs; all other training details remain unchanged (see Section F.2).
Figure 23: Initializing deeper layers of GPT-2 Medium with lazy layers degrades performance. We first train a full GPT-2 Medium model from scratch on OpenWebText for 100K steps. We then construct a second model in which layers 16–24 are initialized using the corresponding lazy layers layers from the trained reference model, while layers 1–15 are randomly initialized. Both models are trained using the same optimization and data settings. The model whose deeper layers are initialized with lazy layers performs significantly worse than its counterpart trained entirely from scratch.
Figure 24: Inheritune continues to provide gains under extended training. Comparison between a full 24-layer GPT-2 Medium model and an 18-layer GPT-2 Medium variant trained using the $\mathsf{Inheritune}$ recipe, both trained on OpenWebText for 200K steps. The $\mathsf{Inheritune}$ model achieves a lower final validation loss (2.86) than the full 24-layer baseline (2.87).
G.5 Discussion on Lazy Layers
In this section, we provide a formal characterization of lazy layers through the rank-based criterion.
Definition G.1 (Lazy Layer (Rank-based)).
Let $\{X_{n}\}_{n=1}^{N}$ be a collection of $N$ sequences, each consisting of $T$ tokens. For each layer $l\in\{1,\dots,L\}$ and head $h\in\{1,\dots,H\}$, let $A_{n,h,l}\in\mathbb{R}^{T\times T}$ denote the attention matrix produced by head $h$ in layer $l$ on input sequence $X_{n}$.
Let $\sigma_{1}(A_{n,h,l})\geq\cdots\geq\sigma_{T}(A_{n,h,l})\geq 0$ be the singular values of $A_{n,h,l}$. For a threshold $\tau\in(0,1)$, we define the approximate rank as
$k^{*}_{n,h,l}(\tau)\;=\;\min\left\{k\in\{1,\dots,T\}\;\middle|\;\frac{\sum_{i=1}^{k}\sigma_{i}(A_{n,h,l})^{2}}{\sum_{j=1}^{T}\sigma_{j}(A_{n,h,l})^{2}}\geq\tau\right\}.$ (2)
We choose $\tau$ via an ablation study; throughout the paper we use $\tau=0.90$ (see Section 4.3).
Define the head-wise aggregated rank and layer-wise aggregated rank as
$\mathrm{Rank}(h,l)\;=\;\frac{1}{N}\sum_{n=1}^{N}k^{*}_{n,h,l}(\tau),\qquad\mathrm{MaxRank}(l)\;=\;\max_{h\in\{1,\dots,H\}}\,\mathrm{Rank}(h,l).$ (3)
For a group of $L$ layers (where $L$ denotes the number of contiguous layers in the group), define
$\mathrm{AvgRank}\;=\;\frac{1}{L}\sum_{l=1}^{L}\mathrm{MaxRank}(l).$ (4)
Lazy Layer. A layer $l$ is a Lazy Layer if $\mathrm{MaxRank}(l)=1$. Lazy Layers. A group of contiguous blocks of $L$ layers is termed Lazy if $\lfloor\mathrm{AvgRank}\rfloor=1$.
Appendix H Theoretical Analysis
H.1 Causal Attention Rank-1 Collapse and Vanishing Gradients
We isolate a simple mechanism in decoder-style (causal) self-attention: when a head becomes highly concentrated on a single sink position across time (e.g., the first/BOS token), the resulting attention matrix is approximately rank-$1$, and the corresponding gradients through the softmax saturate, yielding vanishing updates to $W_{Q}$ and $W_{K}$.
Setup.
We inherit notations discussed in Section 2. Let $X\in\mathbb{R}^{T\times d}$ denote a sequence of $T$ tokens. Define $Q=XW_{Q},\;K=XW_{K},\;V=XW_{V}\in\mathbb{R}^{T\times d}$ and masked logits
$\tilde{A}=\frac{1}{\sqrt{d}}QK^{\top}+M,\qquad M_{ij}=\begin{cases}0&i\geq j,\\ -\infty&i<j,\end{cases}$ (5)